OpenStudy (anonymous):

2 square root of 2 open parentheses square root of 2 minus square root of 10 close parentheses plus square root of 3 open parentheses square root of 3 plus square root of 15

OpenStudy (solomonzelman):

\[2\sqrt{2}(\sqrt{2}-\sqrt{10})+\sqrt{3}(\sqrt{3}+\sqrt{15})\]\[(2\sqrt{2} \times \sqrt{2}) - (2\sqrt{2} \times \sqrt{10})+(\sqrt{3} \times \sqrt{3})+(\sqrt{3} \times \sqrt{15})\]\[(2 \times 4) - (2 \times \sqrt{20}) + (\sqrt{9}) +(\sqrt{45})\]\[8+(2 \times \sqrt{4} \times \sqrt{5}) + 3 + (\sqrt{9} \times \sqrt{5})\]\[8 + 4\sqrt{5} + 3 + 3\sqrt{5}\]add like terms to each other\[8+3+4\sqrt{5}+3\sqrt{5} ⇒11+7\sqrt{5}\]